OpenStudy (anonymous):

Two pipes A and B can separately fill a cistern in 24 and 32 minutes respectively. They started to fill a cistern together, but B is turned off after few minutes and A fills the rest of the cistern. If total time taken is 18 minutes, after how many minutes is tap B turned off?

OpenStudy (anonymous):

first A,B together work next unknow work X by togerther work, add X-18 by b capacity equal to 1 simplify

OpenStudy (perl):

we can use the formula d= r*t. we will be more flexible about d, 1 cistern = (rate to fill up ) * 24 minutes

OpenStudy (perl):

let rA = speed at which pipe A can fill up 1 cistern. , rB = speed at which pipe B can fill up 1 cistern . rA = 1/24 , rB = 1/32 now they work together to fill up 1 cistern, pipe A is fillup the full time of 18 minutes but pipe B is turned off 1 = rA * 18 + rB * t 1 = 1/24 * 18 + 1/32 * t

OpenStudy (perl):

or better yet, rA is the speed at which it fills cisterns

OpenStudy (perl):

ok , if you have any questions let me know :) take care

OpenStudy (perl):

I get t = 8 minutes, so it was turned off after 8 minutes

OpenStudy (perl):

so it is turned off for 10 minutes

OpenStudy (anonymous):

This looks correct.

OpenStudy (anonymous):

So the answer would be: it was turned off after 8 minutes. Good Job @perl